A perfectly competitive market is described by the demand curve QD= 60 – 2P, and the supply curve QS = 5P – 10. A typic...

  1. A perfectly competitive market is described by the demand curve QD= 60 – 2P, and the supply curve QS = 5P – 10. A typical firm has the total cost equation: C = 16 + 2QF + QF2. What is the equilibrium price and quantity in the market? Compute the firm’s total revenue, total cost, and total profit.

MC = dC/dQF = 2QF + 2
Answer #1

For equilibrium in market,

Demand = supply

60-2P = 5P-10

P = 70/7 = $10

Q = 60-2*10 = 40 units

So

Equilibrium price = $10

Equilibrium quantity = 40 units

Total revenue = 10*40 = $400

Total cost = 16 + 2*40 + 40^2

Total cost = $1696

Total profit = 400-1696

Total profit = -$1296